A function is defined as f(x)= 2x+1.the value of X for which f(x), f(2x),f(4x) are in GP
Answer by amarjeeth123(574)

f(x)= 2x+1
f(2x)=4x+1
f(4x)=8x+1
f(x), f(2x),f(4x) are in GP
f(2x)/f(x)=f(4x)/f(2x)
(f(2x))^2=f(x)*f(4x)
(4x+1)^2=(2x+1)(8x+1)
16x^2+8x+1=16x^2+8x+2x+1
2x=0
x=0 is the answer.